### v3

# EXAONE 3.5: Series of Large Language Models for Real-world Use Cases

Abstract
This technical report introduces the EXAONE 3.5 instruction-tuned language models, developed and released by LG AI Research. The EXAONE 3.5 language models are offered in three configurations: 32B, 7.8B, and 2.4B. These models feature several standout capabilities: 1) exceptional instruction following capabilities in real-world scenarios, achieving the highest scores across seven benchmarks, 2) outstanding long-context comprehension, attaining the top performance in four benchmarks, and 3) competitive results compared to state-of-the-art open models of similar sizes across nine general benchmarks. The EXAONE 3.5 language models are open to anyone for research purposes and can be downloaded from https://huggingface.co/LGAI-EXAONE. For commercial use, please reach out to the official contact point of LG AI Research: contact_us@lgresearch.ai.
1 Introduction
EXAONE 3.0 instruction-tuned large language model with 7.8B parameters [research2024exaone3078binstruction] demonstrated strong bilingual capabilities in Korean and English with exceptional real-world performance and instruction-following proficiency. Since its release, we have received diverse feedback from both academic and industrial communities. For instance, academic researchers have emphasized the need for smaller models that can be trained and deployed on low-specification GPUs due to limited access to advanced computational infrastructure. The industry has expressed strong demand for larger models with enhanced performance that remain cost-effective, as well as smaller models suitable for on-device deployment. Additionally, with the increasing adoption of retrieval-augmented generation (RAG) techniques, which generate answers based on reference documents or web search results, there has been substantial demand for models capable of effectively handling longer contexts.
In this report, we present EXAONE 3.5 language models, a collection of instruction-tuned language models ranging from 2.4B to 32B parameters, developed to meet the diverse needs of users. EXAONE 3.5 language models include: 1) 2.4B model optimized for deployment on small or resource-constrained devices, 2) 7.8B model matching the size of its predecessor but offering improved performance, and 3) 32B model delivering exceptional performance. All models support long-context processing of up to 32K tokens. Each model demonstrates state-of-the-art performance in real-world use cases and long-context handling, while remaining competitive in general domains compared to recently released models of similar sizes.
With the release of the EXAONE 3.5 language models, we hope to support researchers to push the boundaries of generative AI and inspire the development of innovative applications that enhance human life. This is in line with the mission of LG AI Research: Advancing AI for a better Life.
2 Model Training
This section describes the detailed information on model configurations and the methods used for pre-training and post-training phases, along with the dataset construction process for each training phase.
2.1 Model Configurations
The EXAONE 3.5 language models are based on the latest decoder-only Transformer architecture, and detailed configurations are described in Table 1. These models are identical in structure to the EXAONE 3.0 7.8B model but mainly differ in their configurations related to sizes. Notably, the EXAONE 3.5 language models extend the maximum context length from 4,096 tokens in EXAONE 3.0 to 32,768 tokens by adopting the long-context fine-tuning [chen2023extendingcontextwindowlarge]. All three models share the same vocabulary, which consists roughly of 50% Korean and 50% English.
Model size 32B 7.8B 2.4B
dd_model 5,120 4,096 2,560
Number of layers 64 32 30
Pre-normalization True True True
Non-linearity SwiGLU [shazeer2020gluvariantsimprovetransformer] SwiGLU SwiGLU
Feedforward dimension 27,392 14,336 7,168
Head type GQA [ainslie2023gqatraininggeneralizedmultiquery] GQA GQA
Number of heads 40 32 32
Number of KV heads 8 8 8
Head size 128 128 80
Max sequence length 32,768 32,768 32,768
RoPE theta [su2023roformerenhancedtransformerrotary] 1,000,000 1,000,000 1,000,000
Tokenizer BBPE [wang2020bbpe] BBPE BBPE
Vocab size 102,400 102,400 102,400
Tied word embedding False False True
Table 1: Configurations of EXAONE 3.5 language models.
2.2 Pre-training
The amount of pre-training corpus data and computational resources are shown in Table 2. The approach to data construction and model training consists of two stages: 1) we perform first-stage pre-training based on the large training corpus, which is collected and processed from as diverse sources as possible aimed to increase the performance on general domains. After that, 2) we collect more data for the domains that need to be strengthened through evaluations and conduct second-stage of pre-training. For instance, we focus on enhancing long-context understanding capabilities in the second-stage.
Model size 32B 7.8B 2.4B
Training tokens 6.5T 9T 6.5T
Amount of computation (FLOPs) $1.25\times 10^{24}$ $4.21\times 10^{23}$ $9.36\times 10^{22}$
Table 2: The sizes of the training data corpus along with the amounts of computation to build EXAONE 3.5 language models.
2.2.1 Context Length Extension
To extend the context length, we utilize the long-context fine-tuning technique [chen2023extendingcontextwindowlarge]. To mitigate the catastrophic forgetting problem [MCCLOSKEY1989109], where the model forgets what it learned during the first pre-training stage, a replay-based method [ai2024yiopenfoundationmodels] is applied. Specifically, during the second-stage pre-training, we reuse a portion of the data used in the first-stage. While documents exceeding the maximum context length is split into smaller chunks in the first-stage, the original corpus are trained without being divided into chunks in the second-stage to extend the models’ context length.
2.2.2 Decontamination
By the nature of massively web-crawled corpus, test-set examples often appear in the training corpus [riddell2024quantifying, xu2024benchmarking]. These contaminated examples are likely to harm generalization performance and confuse test metrics, thus presenting unfair evaluations to users. To prevent the contaminated examples undermine the generalization performance of EXAONE 3.5 language models, we rigorously apply a decontamination process for all targeted benchmark test data and remove contaminated examples from the training pipeline.
We borrow a simple yet powerful substring-level matching method [openai2023GPT4technicalreport] with stricter criteria. The entire decontamination process is described in Figure 4 in Appendix C. We first normalize all test-set examples by removing all other characters except alphabets and numbers, then we extract all unique substrings with sliding window size $S=50$ and a stride of 1. To determine whether a training example is contaminated, we randomly sample $N=10$ substrings from the normalized training example and check if they exist in the substring pools. Table LABEL:tab:decon_example in Appendix C provides examples of documents found in web corpora considered as contaminated.
2.2.3 Training Cost
Considering the computational cost of pre-training a large language model (LLM), it is necessary to make the training efficient by achieving as much high performance as possible with limited resources. Table 3 compares the total amounts of computations required for pre-training between the EXAONE 3.5 32B language model and others of similar size. When we simply approximate the total amounts of computations as the product of the model size and the number of training tokens [kaplan2020scalinglawsneurallanguage, hoffmann2022an], Qwen 2.5 32B, for example, requires 2.77 times more computations than EXAONE 3.5 32B. One of the noticeable characteristics of the EXAONE 3.5 language models is that they demonstrate high performance despite being trained at lower costs than the other baseline models (see Section 3).
Models Model size Training tokens Amount of computation (ratio)
EXAONE 3.5 32B 6.5T 1.00
Qwen 2.5 32B 18T 2.77
Gemma 2 27B 13T 1.69
Yi 1.5 34B 3.6T 0.59
Table 3: Comparison of the total amounts of computations to build models. We approximate the amount of computations as the product of the model size and the number of training tokens. Although the EXAONE 3.5 32B model is behind in the computations compared to Qwen 2.5 and Gemma 2, it has shown competitive performances.
2.3 Post-training
After pre-training, models go through further processes for strengthening their instruction-following capabilities and aligning with human preferences, which are well known as supervised fine-tuning (SFT) [wei2022finetunedlanguagemodelszeroshot] and preference optimization.
2.3.1 Supervised Fine-tuning
To perform well on new or unseen instructions, a model needs to be trained on pairs of instruction-response datasets with varying difficulty from different domains. Hence, in order to build training data covering a wide range of fields, we extract core knowledge from 8M web corpora using a taxonomic system, as shown in Figure 1. We then generate an instruction-tuning dataset based on the extracted knowledge taxonomy. Finally, leveraging an instruction evolution method, which stems from the method proposed in [zeng-etal-2024-automatic], we diversify the complexity levels so that instructions with various complexities and difficulties can be produced.
Figure 1: A procedure of instruction-tuning data construction. First, we extract the core knowledge from large-volume web corpora and classify it within the taxonomy we defined in advance. Next, instruction-tuning data is generated based on the knowledge. To construct additional training data that is more complex, we leverage an instruction-evolving method [zeng-etal-2024-automatic] that lets the final dataset cover various fields with varying levels of difficulty.
2.3.2 Preference Optimization
Direct alignment algorithms (DAAs) [rafailov2024scaling], such as DPO [rafailov2023directpreferenceoptimizationlanguage] and SimPO [meng2024simposimplepreferenceoptimization], are used to train models after supervised fine-tuning to align models with human preferences. We create preference data for training using synthetic data and pre-collected data. For response generation, we sample $N$ responses from multiple models for the prompt $x$ drawn from the preference data and select the best response as $y_{w}$ and the worst response as $y_{l}$ based on the scores of a reward model to create a preference data, $\{x,y_{w},y_{l}\}$. To validate preference data, we use an additional reward model to calculate agreement based on the rankings of the two reward models and filter out data with agreement below the threshold. Our preference optimization comprises multiple stages to sequentially train models $M_{1}$ and $M_{2}$ through DAAs, where $M_{0}$ is initialized from the SFT model. The staged pipeline enables us to mitigate over-optimization [rafailov2024scaling] that may occur during the DAAs’ training process. Figure 2 shows a schematic diagram for constructing our preference dataset and training process.
Figure 2: Overview of the preference optimization pipeline. (Top) Preference Data Creation: It shows the process of constructing preference data $\{x,y_{w},y_{l}\}$ by scoring the responses $y$ generated from a model for the prompt $x$ using a reward model. (Bottom) Preference Optimization: Sequential training process where $M_{0}$ initialized from the SFT model is trained through DAA to obtain $M_{1}$ and $M_{2}$.
2.4 Data Compliance
Developing AI models requires a large amount of data, and the acquisition and utilization of this data can lead to various legal issues, such as copyright infringement, intellectual property infringement, and personal information protection violations. To minimize these risks, LG AI Research conducts AI Compliance reviews throughout the entire process of data collection, AI model training, and information provision. For more detailed information, please refer to the EXAONE 3.0 Technical Report [research2024exaone3078binstruction] and the LG AI Ethics Principles [lgethics].
3 Evaluation
This section presents the evaluation settings and results of EXAONE 3.5 language models on various benchmark datasets. We select recently released open language models for baselines of our models to compare our performances on the benchmarks. All baselines and their detailed information are described in Appendix LABEL:appendix:baseline_details.
3.1 Benchmarks
Considering the diverse nature of user intents, it is crucial for an instruction-tuned model to generate a response aligned to the user’s query, whatever it is. To evaluate our models in comprehensive and various scenarios, we select over a dozen evaluating benchmarks along with a few in-house benchmarks. Table 4 summarizes all benchmarks, which can be grouped into three categories:
  • Real-world Use Cases (Section 3.3): the benchmarks requiring the ability to understand and perform diverse user instructions.
  • Long Context (Section 3.4): the benchmarks evaluating the ability to understand the long context.
  • General Domain (Section 3.5): the benchmarks embracing general domain abilities that LLMs are expected to have. Specifically, this category includes benchmarks for measuring the ability to solve mathematical problems, the ability to write source codes, and the parametric knowledge embedded in an LLM.
Category Benchmark Lang Evaluation Settings Metric
Real-world Use Cases MT-Bench [zheng2023judging] EN LLM-as-a-judge (judge: gpt-4o-2024-08-06) [^0] The separability of the original GPT-4 judge results is notably low, prompting the adoption of gpt-4o-2024-08-06 as judge. LLM score
LiveBench [white2024livebenchchallengingcontaminationfreellm] (v2024-08-31) EN Ground-truth match Accuracy
Arena-Hard-v0.1 [li2024crowdsourceddatahighqualitybenchmarks] EN LLM-as-a-judge (judge: gpt-4-1106-preview) Win rate
AlpacaEval 2.0 LC [dubois2024lengthcontrolledalpacaevalsimpleway] EN LLM-as-a-judge (judge: gpt-4-1106-preview) Win rate
IFEval[zhou2023instructionfollowingevaluationlargelanguage] EN Prompt-level / strict accuracy Accuracy
KoMT-Bench [komt-bench] KO LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
LogicKor [logickor] KO LLM-as-a-judge (judge: gpt-4-1106-preview) LLM score
Long Context Needle-In-A-Haystack [kamradt2023llmtest] EN/KO Ground-truth match Accuracy
LongBench [bai-etal-2024-longbench] EN Ground-truth match F1, Rouge
LongRAG [jiang2024longragenhancingretrievalaugmentedgeneration] (extended) EN LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
Ko-LongRAG (In-house) KO LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
Ko-WebRAG (In-house) KO LLM-as-a-judge (judge: gpt-4o-2024-08-06) LLM score
General Domain GSM8K [cobbe2021trainingverifierssolvemath] EN 0-shot / CoT Accuracy
MATH [hendrycks2021measuringmathematicalproblemsolving, minerva_math] EN 0-shot / CoT Accuracy
HumanEval [chen2021evaluatinglargelanguagemodels] EN 0-shot pass@1
MBPP [austin2021programsynthesislargelanguage] EN 0-shot (Evalplus base) [^1] We choose the MBPP base from EvalPlus [liu2023is], which is a subset of the original and consists of refined, high-quality problems. pass@1
GPQA [rein2023gpqagraduatelevelgoogleproofqa] EN 0-shot / CoT Accuracy
ARC-C [clark2018thinksolvedquestionanswering] EN 0-shot Accuracy
BBH [suzgun2022challenging] EN 0-shot / CoT Accuracy
MMLU [hendrycks2020measuring] EN 0-shot / CoT Accuracy
KMMLU [son2024kmmlumeasuringmassivemultitask] KO 0-shot / CoT Accuracy
Table 4: The benchmarks used to evaluate the performance of EXAONE 3.5 language models along with their target languages, evaluation settings, and the metrics. LongRAG is extended from the original, and Ko-LongRAG and Ko-WebRAG are in-house benchmarks (see Section 3.4).
3.2 Overall Performance
The results of overall performance against three categories are presented in Table 5. Our EXAONE 3.5 language models, with sizes 32B and 7.8B, perform best in Real-world Use Cases and Long Context categories compared to baseline models while showing competitive results in the General Domain category. Our smallest model, EXAONE 3.5 2.4B, outperforms baselines with similar sizes in all three categories, demonstrating strong performance. Surprisingly, our 2.4B model, despite its small size, has shown better performance compared to baselines even with a larger size (< 9B) except for Qwen 2.5 7B in General Domain. Considering the recent surge in demand for smaller large language models (sLLM) [sllmsurvey2024], we believe that our EXAONE 3.5 2.4B model is well-positioned to be highly competitive in both academic and industrial use.
In the following sections, we elaborate on detailed evaluation settings and the results for each category.
Models Real-world Use Cases Long Context General Domain
EXAONE 3.5 32B 74.3 71.1 74.8
Qwen 2.5 32B [qwen2.5] 69.8 66.9 78.7
C4AI Command R 32B  [cohere_for_ai_2024] 46.0 63.4 56.8
Gemma 2 27B [gemmateam2024gemma2improvingopen] 64.2 - 68.7
Yi 1.5 34B [ai2024yiopenfoundationmodels] 46.9 - 53.9
EXAONE 3.5 7.8B 70.7 66.6 70.2
Qwen 2.5 7B [qwen2.5] 52.7 56.1 71.0
Llama 3.1 8B [grattafiori2024llama3herdmodels] 48.6 58.8 62.4
Gemma 2 9B [gemmateam2024gemma2improvingopen] 57.9 - 62.9
Phi 3 small (7B) [abdin2024phi3technicalreporthighly] 41.7 33.4 63.2
EXAONE 3.5 2.4B 61.1 63.4 63.3
Qwen 2.5 3B [qwen2.5] 44.5 40.7 62.1
Qwen 2.5 1.5B [qwen2.5] 30.1 34.5 47.9
Llama 3.2 3B [llama3.2] 36.7 44.2 54.9
Gemma 2 2B [gemmateam2024gemma2improvingopen] 41.7 - 42.2
Table 5: Overall comparison results of EXAONE 3.5 language models with similar-sized baseline language models. Here, a dash (-) indicates the model does not support context lengths longer than 16K. Bold scores indicate the best performance, and underlined scores mean the second best. The detailed information for each baseline is described in Appendix LABEL:appendix:baseline_details.
3.3 Real-world Use Cases
For the Real-world Use Cases category, we have compiled seven benchmarks that represent real-world queries users might submit to a chatbot model. In MT-Bench, KoMT-Bench, and LogicKor, models’ responses consisting of multi-turns are evaluated by a judge model. For Arena-Hard and AlpacaEval, responses of a target language model are compared with those of a reference model (gpt-4-0314 and gpt-4-1106-preview, respectively) by a judge model, recording the win rate. LiveBench (ver. 2024-08-31) and IFEval (prompt-strict) assess how well the models’ responses align with user instructions by matching them to the ground-truth responses.
Models MT-Bench LiveBench Arena-Hard AlpacaEval IFEval KoMT-Bench LogicKor Average
EXAONE 3.5 32B 8.51 43.0 78.6 60.6 81.7 8.05 9.06 74.3
Qwen 2.5 32B 8.49 50.6 67.0 41.0 78.7 7.75 8.89 69.8
C4AI Command R 32B 7.38 29.7 17.0 25.9 26.1 6.72 8.24 46.0
Gemma 2 27B 8.28 40.0 57.5 52.2 59.7 7.19 8.56 64.2
Yi 1.5 34B 7.64 26.2 23.1 34.8 55.5 4.88 6.33 46.9
EXAONE 3.5 7.8B 8.29 39.8 68.7 54.2 78.9 7.96 9.08 70.7
Qwen 2.5 7B 6.48 35.6 48.9 31.7 72.5 5.19 6.38 52.7
Llama 3.1 8B 7.59 28.3 27.7 25.7 74.5 4.85 5.99 48.6
Gemma 2 9B 7.64 32.1 43.6 47.3 54.7 7.10 8.05 57.9
Phi 3 small (7B) 7.63 27.9 26.8 29.2 59.5 3.22 3.99 41.7
EXAONE 3.5 2.4B 7.81 33.0 48.2 37.1 73.6 7.24 8.51 61.1
Qwen 2.5 3B 7.21 25.7 26.4 17.4 60.8 5.68 5.21 44.5
Qwen 2.5 1.5B 5.72 19.2 10.6 8.4 40.7 3.87 3.60 30.1
Llama 3.2 3B 6.94 24.0 14.2 18.7 70.1 3.16 2.86 36.7
Gemma 2 2B 7.20 20.0 19.1 29.1 50.5 4.83 5.29 41.7
Table 6: Performance comparison results of EXAONE 3.5 language models with similar-sized recently-released language models on seven benchmarks representing real-world use case scenarios. When calculating the macro average, the scores of MT-Bench, KoMT-Bench, and LogicKor are multiplied by 10 because they are scored out of 10 and the rest are scored out of 100. Bold scores indicate the best performance, and underlined scores mean the second best.
As presented in Table 6, our three models have shown the best performance against baselines of similar size in all benchmarks, except for the 32B model in LiveBench. Furthermore, by outperforming others in both English and Korean benchmarks, EXAONE 3.5 language models demonstrate their superior bilingual abilities.
3.4 Long Context
The ability to process and understand long contexts is increasingly important for modern LLMs, as it enables their application in more complex scenarios. To demonstrate EXAONE 3.5 language models’ long context performance, we evaluate our models using benchmarks designed for a synthetic task with long context inputs, along with various retrieval-augmented generation (RAG) benchmarks.
3.4.1 Needle-in-a-Haystack
Needle-in-a-Haystack (NIAH) [kamradt2023llmtest] serves as a benchmark to assess how effectively models can locate and retrieve information hidden at random locations within long documents. We comprehensively evaluate our models’ ability to process and retrieve information from long contexts, up to 32K tokens. Furthermore, we extend NIAH to Korean and employ it to evaluate our models’ long context processing ability across both English and Korean contexts.
Figure 3 demonstrates that our models achieve near-perfect accuracy in retrieving targeted information across all tested document depths and context lengths in both English and Korean. These results highlight their robust long context processing capabilities, particularly in tasks demanding precise information retrieval and complex reasoning.
Figure 3: NIAH results of EXAONE 3.5 language models. The x-axis represents the token length of the input text, while the y-axis shows the relative position within the text, expressed as a percentage (0% corresponds to the beginning, and 100% to the end). The results are represented using a color-coded scheme: green indicates successful retrievals, and red represents unsuccessful ones. EXAONE 3.5 language models achieve near-perfect accuracy in retrieving information across various document depths and context lengths in English and Korean.
3.4.2 Long Context Understanding
To assess long context understanding capabilities, we evaluate our models using benchmarks including LongBench [bai-etal-2024-longbench] and LongRAG [jiang2024longragenhancingretrievalaugmentedgeneration]. We expand unanswerable cases in LongRAG to make it more challenging. We also build Ko-LongRAG, the Korean counterpart to LongRAG, to evaluate long context understanding in Korean. For more realistic RAG scenario, requiring answers to difficult questions using actual web-searched results, we constructed Ko-WebRAG benchmark. We refer readers to the Appendix LABEL:appendix:lc_setting for more details.
As shown in Table 7, EXAONE 3.5 language models have shown superior performance compared to other models [^2] Gemma models and Yi 1.5 34B model are excluded from evaluations due to their context limits ($\leq$ 16k tokens), ensuring fair comparison., except for the 32B and 7.8B models in LongBench. When averaged across the benchmarks, our three models outperform all baselines, confirming their capabilities to process complex, extended contexts effectively.
Models LongBench LongRAG Ko-LongRAG Ko-WebRAG Average
EXAONE 3.5 32B 49.2 67.6 85.3 82.3 71.1
Qwen 2.5 32B 49.1 63.6 73.5 81.3 66.9
C4AI Command R 32B 50.9 55.3 72.3 75.0 63.4
Gemma 2 27B - - - - -
Yi 1.5 34B - - - - -
EXAONE 3.5 7.8B 46.0 68.3 71.7 80.3 66.6
Qwen 2.5 7B 47.2 60.1 55.3 61.7 56.1
Llama 3.1 8B 44.6 55.1 64.8 70.7 58.8
Gemma 2 9B - - - - -
Phi 3 small (7B) 40.6 52.7 7.7 32.7 33.4
EXAONE 3.5 2.4B 42.7 63.3 74.7 73.0 63.4
Qwen 2.5 3B 42.0 45.8 40.5 34.7 40.7
Qwen 2.5 1.5B 37.1 39.0 33.8 28.0 34.5
Llama 3.2 3B 41.7 45.9 39.3 50.0 44.2
Gemma 2 2B - - - - -
Table 7: Performance comparison results of EXAONE 3.5 language models with similar-sized recently released language models across four benchmarks representing long context scenarios. A dash (-) indicates that the model does not support context lengths longer than 16K. Context lengths for each model are detailed in Table LABEL:tab:baselines. The average score in the rightmost is calculated as a macro average across the benchmarks. Bold scores indicate the best performance, and underlined scores mean the second best.
3.5 General Domain
Language models are now expected to achieve human-level capabilities in various general domains, such as solving mathematical problems or writing source code programs. To evaluate overall performance in the general domains, we select nine benchmarks in three main domains: 1) GSM8K (CoT) and MATH (CoT) for mathematics, 2) HumanEval (Evalplus base) and MBPP (Evalplus base) for coding, and 3) MMLU (CoT), KMMLU (CoT), GPQA (CoT), ARC-C, and BBH (CoT) for assessing the amount of knowledge embedded in an LLM.
To better simulate the real-world scenarios where a chatbot model usually receives a single query from users, we evaluate all benchmarks in the General Domain category using the 0-shot setting. To achieve this, we prompt language models with instructions that require specific answer formats and parse the final answer from the responses. For a fair comparison, we use the same prompts across all models. We make public all the prompts we used in Appendix LABEL:appendix:general_dom_setting for transparent reproducibility.
Models GSM8K MATH HumanEval MBPP MMLU KMMLU GPQA ARC-C BBH Average
EXAONE 3.5 32B 91.9 70.5 87.2 81.8 78.3 57.0 39.7 91.7 75.3 74.8
Qwen 2.5 32B 92.0 76.5 89.0 88.9 81.4 62.1 40.9 95.1 82.7 78.7
C4AI Command R 32B 56.5 24.3 68.3 78.8 71.1 41.5 27.4 88.0 55.7 56.8
Gemma 2 27B 84.2 49.4 79.3 80.7 74.8 53.8 33.6 92.9 69.7 68.7
Yi 1.5 34B 83.7 52.0 5.5 35.7 75.3 41.7 30.0 93.9 67.6 53.9
EXAONE 3.5 7.8B 87.6 69.8 84.2 79.4 69.0 52.4 32.5 87.6 69.7 70.2
Qwen 2.5 7B 90.4 70.4 82.3 78.8 73.1 49.9 33.1 90.6 70.1 71.0
Llama 3.1 8B 82.1 48.8 67.7 70.6 72.4 45.9 27.4 83.7 63.3 62.4
Gemma 2 9B 82.0 44.6 68.3 75.1 73.7 34.6 27.9 90.5 69.7 62.9
Phi 3 small (7B) 86.3 47.8 72.6 72.0 68.8 33.4 25.3 90.4 72.5 63.2
EXAONE 3.5 2.4B 82.5 60.2 76.2 74.3 60.4 45.8 28.4 79.2 62.9 63.3
Qwen 2.5 3B 84.3 61.4 72.6 72.5 61.0 41.7 25.8 82.1 57.3 62.1
Qwen 2.5 1.5B 69.8 48.5 55.5 65.6 48.8 5.0 23.1 72.4 42.2 47.9
Llama 3.2 3B 77.4 46.6 54.9 60.6 64.9 35.0 23.2 78.0 53.8 54.9
Gemma 2 2B 29.8 18.7 45.7 55.0 56.1 37.4 22.6 76.3 38.2 42.2
Table 8: Performance comparison results of EXAONE 3.5 models with similar-sized recently-released language models on nine benchmarks representing general scenarios. The macro average is used to evaluate the overall performance. Bold scores indicate the best performance, and underlined scores mean the second best.
Table 8 shows the results of EXAONE 3.5 language models and their baseline models on the benchmarks in the General Domain category. When averaged across the benchmarks, our EXAONE 3.5 language models with sizes 32B and 7.8B demonstrate competitive performance compared to baselines of similar size. The EXAONE 3.5 2.4B model, on the other hand, outperforms all baselines in the average score.
4 Responsible AI
EXAONE 3.5 language models were developed in accordance with the Responsible AI Development Framework encompassing data governance, ethical considerations, and risk management as it would be made available to a wide range of users. Given the nature of open models – eventually leading to wide use in various domains – we aim to maximize social benefits while ensuring humanity, fairness, safety, accountability, and transparency as mandated by the LG AI Ethics Principles [lgethics].
4.1 Benefits
EXAONE 3.5 language models are open for research purposes, aiming to advance AI research. Based on the feedback we have received since the release of the EXAONE 3.0 7.8B model, we now offer models of more diverse sizes: 2.4B, 7.8B, and 32B. This will allow researchers to select an optimal model for their research objectives and computing environment. We hope that this flexibility will support a wide spectrum, ranging from foundational research to domain-specific applications. It is also expected to contribute positively to the advancement of generative AI, building upon the significant performance improvements over previous version.
To ensure the reliability of the release, we have implemented a standardized data compliance protocol, guaranteeing high-quality data. This standardized approach provides a trustworthy foundation for researchers to use the model across various research areas in the future.
While external users can employ EXAONE 3.5 language models in diverse domains, precisely identifying specific user needs has been challenging. To address this, we have conducted extensive reviews of its applicability across a wide range of domains. Additionally, we have collaborated closely with LG affiliates, including business and research teams, to better align the model with specific user requirements.
4.2 Risks and Mitigations
Open models can positively contribute to the AI community, but there are challenges in ensuring responsible use. We conducted an AI ethical impact assessment to identify potential risks such as unintended inequality and discrimination against socially disadvantaged groups, the generation of harmful content, and malicious misuse by users. We have adopted various policies and research initiatives to mitigate the potential risks identified through this assessment.
First, on the data side, we conducted a legal risk assessment on all candidate datasets to enhance privacy and security. Based on the outcomes, we determined the suitability of each dataset for training and performed a de-identification process to remove sensitive data from qualified dataset. To minimize bias in the training data and ensure data quality, we documented all pre-processing steps and adopted a standardized data processing protocol. Considering practical difficulties of verifying the representativeness of all data, we conducted a qualitative evaluation of a small sample of data. For a quantitative evaluation, we endeavored to minimize data-related risks by verifying the data subsets through performance evaluation after the model training was completed. Also, we carefully reviewed the open-source libraries used in our model development.
The levels of AI ethical considerations and regulatory requirements may vary across different user needs and characteristics (e.g., country of residence, age, etc.). To address this, we will continue to monitor global AI regulations and take immediate action as needed to avoid potential regulatory violations. A lack of transparency in an AI model’s decision-making process can reduce trust among users and stakeholders. To address this limitation, we continuously analyze and evaluate our model’s performance to identify weaknesses and areas for improvement. While fully explaining AI model’s decision-making process remains challenging, we are committed to advancing explainability through ongoing research.
4.3 Safety
We conducted comprehensive evaluations of EXAONE 3.5 language models’ ethics and security using a third-party dataset: Korean Large Language Model Trustworthiness Benchmark Data [NIARedteaming], provided by the Ministry of Science and ICT of the Republic of Korea and the National Information Society Agency (NIA). This dataset is specifically designed to assess the harmlessness of language models. The evaluation results are presented in Table 9. To measure the performance, we asked a model to choose one of five options. If the selected option is included in the set of correct answers, then it is scored as correct. In the provided dataset, the first two options were labeled “False” and the remaining three were labeled “True”. To mitigate potential bias from the order of options, we shuffled the order of options randomly for each evaluation. While the experimental results demonstrated effectiveness in filtering harmful reactions, there is still room for improvement.
Category Subcategory Test Cases Accuracy
32B 7.8B 2.4B
Bias Gender & Sexual Orientation 295 91.2% 87.5% 76.6%
Race & Ethnicity & Nationality 432 86.8% 85.0% 72.2%
Political Affiliation 720 82.8% 79.9% 56.7%
Region 415 87.7% 84.6% 69.2%
Job 442 86.2% 81.9% 67.0%
Miscellaneous 406 85.2% 86.5% 73.2%
Subtotal 2,710 86.0% 83.5% 67.4%
Hate Gender & Sexual Orientation 399 95.2% 92.2% 83.5%
Race & Ethnicity & Nationality 749 91.6% 88.4% 73.8%
Political Affiliation 1,164 85.7% 83.4% 66.2%
Region 499 92.0% 87.2% 74.1%
Job 852 91.0% 87.8% 72.3%
Subtotal 3,663 90.0% 86.9% 72.2%
Illegal Illegal 1,126 92.9% 89.6% 80.3%
Sensitiveness Contentious 710 83.1% 86.1% 79.0%
Ethical 966 81.2% 83.7% 72.8%
Predictive 825 79.8% 82.3% 71.0%
Subtotal 2,501 81.2% 83.9% 74.0%
Overall 10,000 87.1% 85.6% 72.2%
Table 9: Evaluation results of EXAONE 3.5 language models on the Korean Large Language Model Trustworthiness Benchmark Data [NIARedteaming] to assess the model’s harmlessness. The accuracy is determined by the number of times the model selects appropriate options when presented with questions involving various harmful and dangerous categories, such as illegal content.
5 Limitations
EXAONE 3.5 language models, like all existing language models, have certain limitations and may occasionally generate inappropriate responses. The language model generates responses based on the output probability of tokens, and it is determined during learning from training data. While we have made every effort to exclude personal, harmful, and biased information from the training data, some problematic content may still be included, potentially leading to undesirable responses. Please note that the text generated by EXAONE 3.5 language models does not reflect the views of LG AI Research.
  • Inappropriate answers may be generated, which contain personal, harmful or other inappropriate information.
  • Biased responses may be generated, which are associated with age, gender, race, and so on.
  • The generated responses rely heavily on statistics from the training data, which can result in the generation of semantically or syntactically incorrect sentences.
  • Since the models do not reflect the latest information, the responses may be false or contradictory.
LG AI Research strives to reduce potential risks that may arise from EXAONE 3.5 language models. Users are not allowed to engage in any malicious activities (e.g., keying in illegal information) that may induce the creation of inappropriate outputs violating LG AI’s ethical principles when using EXAONE 3.5 language models.
6 Deployment
Section B in Appendix provides license information for using the EXAONE 3.5 language models. Understanding the license information is essential for the legal utilization of the language model.
7 Conclusion
In response to the growing interest from academia and industry, we are excited to release EXAONE 3.5 language models that excel in real-world use cases and long-context understanding. These models are available in three sizes (32B, 7.8B, and 2.4B).
To validate performance of our models in the real-world use case scenarios, we evaluated our models on seven benchmarks requiring diverse instructions understanding. To assess long-context understanding, we evaluated our models on four benchmarks. Our models consistently outperformed in both categories. Additionally, our models exhibited competitive performance in general domains including solving mathematical problems and writing code. In particular, our 2.4B model ranked first in average scores across general domains.
Our models are available to everyone for research purposes, and we welcome your feedback to help us improve the models. If you have any feedback or are interested in exploring commercial opportunities with our models, please reach out to contact_us@lgresearch.ai.
Appendix A Contributors
All authors are listed in alphabetical order by last name.
Core Contributors
Eunbi Choi, Kibong Choi, Seokhee Hong, Junwon Hwang, Hyojin Jeon, Hyunjik Jo, Joonkee Kim, Seonghwan Kim, Soyeon Kim, Sunkyoung Kim, Yireun Kim, Yongil Kim, Haeju Lee, Jinsik Lee, Kyungmin Lee, Sangha Park, Heuiyeen Yeen, Hyeongu Yun
Contributors
Soyoung An, Kyunghoon Bae, Stanley Jungkyu Choi, Gerrard Jeongwon Jo, Jiyeon Jung, Yountae Jung, Hyosang Kim, Youchul Kim, Edward Hwayoung Lee, Honglak Lee, Woohyung Lim, Sooyoun Park, Yongmin Park, Sihoon Yang
Appendix B Model License
EXAONE AI Model License Agreement 1.1 - NC
This License Agreement (“Agreement”) is entered into between you (“Licensee”) and LG Management Development Institute Co., Ltd. (“Licensor”), governing the use of the EXAONE AI Model (“Model”). By downloading, installing, copying, or using the Model, you agree to comply with and be bound by the terms of this Agreement. If you do not agree to all the terms, you must not download, install, copy, or use the Model. This Agreement constitutes a binding legal agreement between the Licensee and Licensor.
1. Definitions
1.1 Model: The artificial intelligence model provided by Licensor, which includes any software, algorithms, machine learning models, or related components supplied by Licensor. This definition extends to encompass all updates, enhancements, improvements, bug fixes, patches, or other modifications that may be provided by Licensor from time to time, whether automatically or manually implemented.
1.2 Derivatives: Any modifications, alterations, enhancements, improvements, adaptations, or derivative works of the Model created by Licensee or any third party. This includes changes made to the Model’s architecture, parameters, data processing methods, or any other aspect of the Model that results in a modification of its functionality or output.
1.3 Output: Any data, results, content, predictions, analyses, insights, or other materials generated by the Model or Derivatives, regardless of whether they are in their original form or have been further processed or modified by the Licensee. This includes, but is not limited to, textual or numerical produced directly or indirectly through the use of the Model.
1.4 Licensor: LG Management Development Institute Co., Ltd., the owner, developer, and provider of the EXAONE AI Model. The Licensor holds all rights, title, and interest in the Model and is responsible for granting licenses to use the Model under the terms specified in this Agreement.
1.5 Licensee: The individual, organization, corporation, academic institution, government agency, or other entity using or intending to use the Model under the terms and conditions of this Agreement. The Licensee is responsible for ensuring compliance with the Agreement by all authorized users who access or utilize the Model on behalf of the Licensee.
2. License Grant
2.1 Grant of License: Subject to the terms and conditions outlined in this Agreement, the Licensor hereby grants the Licensee a limited, non-exclusive, non-transferable, worldwide, and revocable license to:
a. Access, download, install, and use the Model solely for research purposes. This includes evaluation, testing, academic research, experimentation, and participation in competitions, provided that such participation is in a non-commercial context. Notwithstanding Section 3.1, the Licensee may only provide the Model or Derivatives for a competition if no commercial license is granted to the competition organizer or any third party.
b. Publicly disclose research results and findings derived from the use of the Model or Derivatives, including publishing papers or presentations.
c. Modify the Model and create Derivatives based on the Model, provided that such modifications and Derivatives are used exclusively for research purposes. The Licensee may conduct experiments, perform analyses, and apply custom modifications to the Model to explore its capabilities and performance under various scenarios. If the Model is modified, the modified Model must include “EXAONE” at the beginning of its name.
d. Distribute the Model and Derivatives in each case with a copy of this Agreement.
2.2 Scope of License: The license granted herein does not authorize the Licensee to use the Model for any purpose not explicitly permitted under this Agreement. Any use beyond the scope of this license, including any commercial application or external distribution, is strictly prohibited unless explicitly agreed upon in writing by the Licensor.
3. Restrictions
3.1 Commercial Use: The Licensee is expressly prohibited from using the Model, Derivatives, or Output for any commercial purposes, including but not limited to, developing or deploying products, services, or applications that generate revenue, whether directly or indirectly. Any commercial exploitation of the Model or its derivatives requires a separate commercial license agreement with the Licensor. Furthermore, the Licensee shall not use the Model, Derivatives or Output to develop or improve other models.
3.2 Reverse Engineering: The Licensee shall not decompile, disassemble, reverse engineer, or attempt to derive the source code, underlying ideas, algorithms, or structure of the Model, except to the extent that such activities are expressly permitted by applicable law. Any attempt to bypass or circumvent technological protection measures applied to the Model is strictly prohibited.
3.3 Unlawful Use: The Licensee shall not use the Model and Derivatives for any illegal, fraudulent, or unauthorized activities, nor for any purpose that violates applicable laws or regulations. This includes but is not limited to the creation, distribution, or dissemination of malicious, deceptive, or unlawful content.
3.4 Ethical Use: The Licensee shall ensure that the Model or Derivatives is used in an ethical and responsible manner, adhering to the following guidelines:
a. The Model and Derivatives shall not be used to generate, propagate, or amplify false, misleading, or harmful information, including fake news, misinformation, or disinformation.
b. The Model and Derivatives shall not be employed to create, distribute, or promote content that is discriminatory, harassing, defamatory, abusive, or otherwise offensive to individuals or groups based on race, gender, sexual orientation, religion, nationality, or other protected characteristics.
c. The Model and Derivatives shall not infringe on the rights of others, including intellectual property rights, privacy rights, or any other rights recognized by law. The Licensee shall obtain all necessary permissions and consents before using the Model and Derivatives in a manner that may impact the rights of third parties.
d. The Model and Derivatives shall not be used in a way that causes harm, whether physical, mental, emotional, or financial, to individuals, organizations, or communities. The Licensee shall take all reasonable measures to prevent misuse or abuse of the Model and Derivatives that could result in harm or injury.
4. Ownership
4.1 Intellectual Property: All rights, title, and interest in and to the Model, including any modifications, Derivatives, and associated documentation, are and shall remain the exclusive property of the Licensor. The Licensee acknowledges that this Agreement does not transfer any ownership rights to the Licensee. All trademarks, service marks, and logos associated with the Model are the property of the Licensor.
4.2 Output: All rights, title, and interest in and to the Output generated by the Model and Derivatives whether in its original form or modified, are and shall remain the exclusive property of the Licensor. Licensee may use, modify, and distribute the Output and its derivatives for research purpose. The Licensee shall not claim ownership of the Output except as expressly provided in this Agreement. The Licensee may use the Output solely for the purposes permitted under this Agreement and shall not exploit the Output for unauthorized or commercial purposes.
4.3 Attribution: In any publication or presentation of results obtained using the Model, the Licensee shall provide appropriate attribution to the Licensor, citing the Model’s name and version, along with any relevant documentation or references specified by the Licensor.
5. No Warranty
5.1 “As-Is” Basis: The Model, Derivatives, and Output are provided on an “as-is” and “as-available” basis, without any warranties or representations of any kind, whether express, implied, or statutory. The Licensor disclaims all warranties, including but not limited to, implied warranties of merchantability, fitness for a particular purpose, accuracy, reliability, non-infringement, or any warranty arising from the course of dealing or usage of trade.
5.2 Performance and Reliability: The Licensor does not warrant or guarantee that the Model, Derivatives or Output will meet the Licensee’s requirements, that the operation of the Model, Derivatives or Output will be uninterrupted or error-free, or that defects in the Model will be corrected. The Licensee acknowledges that the use of the Model, Derivatives or Output is at its own risk and that the Model, Derivatives or Output may contain bugs, errors, or other limitations.
5.3 No Endorsement: The Licensor does not endorse, approve, or certify any results, conclusions, or recommendations derived from the use of the Model. The Licensee is solely responsible for evaluating the accuracy, reliability, and suitability of the Model for its intended purposes.
6. Limitation of Liability
6.1 No Liability for Damages: To the fullest extent permitted by applicable law, in no event shall the Licensor be liable for any special, incidental, indirect, consequential, exemplary, or punitive damages, including but not limited to, damages for loss of business profits, business interruption, loss of business information, loss of data, or any other pecuniary or non-pecuniary loss arising out of or in connection with the use or inability to use the Model, Derivatives or any Output, even if the Licensor has been advised of the possibility of such damages.
6.2 Indemnification: The Licensee agrees to indemnify, defend, and hold harmless the Licensor, its affiliates, officers, directors, employees, and agents from and against any claims, liabilities, damages, losses, costs, or expenses (including reasonable attorneys’ fees) arising out of or related to the Licensee’s use of the Model, any Derivatives, or any Output, including any violation of this Agreement or applicable laws.
7. Termination
7.1 Termination by Licensor: The Licensor reserves the right to terminate this Agreement and revoke the Licensee’s rights to use the Model at any time, with or without cause, and without prior notice if the Licensee breaches any of the terms or conditions of this Agreement. Termination shall be effective immediately upon notice.
7.2 Effect of Termination: Upon termination of this Agreement, the Licensee must immediately cease all use of the Model, Derivatives, and Output and destroy all copies of the Model, Derivatives, and Output in its possession or control, including any backup or archival copies. The Licensee shall certify in writing to the Licensor that such destruction has been completed.
7.3 Survival: The provisions of this Agreement that by their nature should survive termination, including but not limited to, Sections 4 (Ownership), 5 (No Warranty), 6 (Limitation of Liability), and this Section 7 (Termination), shall continue to apply after termination.
8. Governing Law
8.1 Governing Law: This Agreement shall be governed by and construed in accordance with the laws of the Republic of Korea, without regard to its conflict of laws principles.
8.2 Arbitration: Any disputes, controversies, or claims arising out of or relating to this Agreement, including its existence, validity, interpretation, performance, breach, or termination, shall be referred to and finally resolved by arbitration administered by the Korean Commercial Arbitration Board (KCAB) in accordance with the International Arbitration Rules of the Korean Commercial Arbitration Board in force at the time of the commencement of the arbitration. The seat of arbitration shall be Seoul, Republic of Korea. The tribunal shall consist of one arbitrator. The language of the arbitration shall be English.
9. Alterations
9.1 Modifications: The Licensor reserves the right to modify or amend this Agreement at any time, in its sole discretion. Any modifications will be effective upon posting the updated Agreement on the Licensor’s website or through other means of communication. The Licensee is responsible for reviewing the Agreement periodically for changes. Continued use of the Model after any modifications have been made constitutes acceptance of the revised Agreement.
9.2 Entire Agreement: This Agreement constitutes the entire agreement between the Licensee and Licensor concerning the subject matter hereof and supersedes all prior or contemporaneous oral or written agreements, representations, or understandings. Any terms or conditions of any purchase order or other document submitted by the Licensee in connection with the Model that are in addition to, different from, or inconsistent with the terms and conditions of this Agreement are not binding on the Licensor and are void.
By downloading, installing, or using the EXAONE AI Model, the Licensee acknowledges that it has read, understood, and agrees to be bound by the terms and conditions of this Agreement.
Appendix C Decontamination Details
As described in Section 2.2.2, we apply the decontamination process over our training data to remove any data instances that overlap with test sets, thus harming the generalization performance of our models. Figure 4 presents an overview of our decontamination process, and Table LABEL:tab:decon_example shows examples of contaminated, therefore removed data.
Figure 4: A summary of the decontamination method employed to train EXAONE 3.5 language models. Adopting an approach borrowed from the GPT-4 method, we increase the number of random sample to $N=10$ for stricter decontamination.
Benchmark Benchmark example Contaminated web corpus
MMLU [hendrycks2020measuring] A teacher has three packages of stickers. One package contains 56 stickers, another package contains 48 stickers, and the third package contains 58 stickers. If the teacher divides all the stickers equally among 27 students, how many stickers will each student receive?
A. 6 stickers
B. 9 stickers
C. 54 stickers
D. 81 stickers
Answer:
(…truncated…) A teacher has three packages of stickers. One package contains 56 stickers, another package contains 48 stickers, and the third package contains 58 stickers. If the teacher divides all the stickers equally among 27 students, how many stickers will each student receive?
6 stickers is correct
#4 Last week Mario walked 7 3/4 miles. This week he walked 15 5/6 miles. What is the difference between the distance he walked this week and the distance he walked last week? (…truncated…)
KMMLU [son2024kmmlumeasuringmassivemultitask]